Jump to main content.


Drake Chemical

EPA Daily Situation Reports for December 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Thursday, December 31, 1998

FOR DATES: From 7:00 a.m. Wednesday, Dec. 30, 1998 to 7:00 a.m. Thursday, Dec. 31, 1998.

WEATHER: Partly cloudy and windy. The temperature range was 15oF to 23oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 760.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, drum handling, feed preparation, lime mixing, debris separation, debris handling, wood shredding, ash handling and ash backfilling. Excavation activities were continued in the vicinity of the former debris separation building. Several intact drums and several drum carcasses were unearthed during excavation activities. The recently excavated drums were sampled and overpacked on this date. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-12-26-98-315) for bin 2. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, December 30, 1998

FOR DATES: From 7:00 a.m. Tuesday, Dec. 29, 1998 to 7:00 a.m. Wednesday, Dec. 30, 1998.

WEATHER: Overcast w/ light rain and snow showers. The temperature range was 26oF to 33oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 695.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.1 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, drum handling, feed preparation, debris separation, debris handling, wood shredding, ash handling and ash backfilling. Excavation activities were continued in the vicinity of the former debris separation building. Several intact drums and several drum carcasses were unearthed during excavation activities. The weekly NPDES samples were collected from the WWTP effluent. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-12-24-98-313 and bot-ash-12-25-98) for bin 4 and bin 3, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-12-25-98-063) for bin B. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling, feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs and Larry Newcomer (PADEP).

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Tuesday, December 29, 1998

FOR DATES: From 7:00 a.m. Monday, Dec. 28, 1998 to 7:00 a.m. Tuesday, Dec. 29, 1998. WEATHER: Overcast. The temperature range was 23oF to 35oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 728.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.7 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris separation, debris handling, rock backfilling, ash handling and ash backfilling. Excavation activities were continued in the vicinity of the former debris separation building. Several intact drums and several drum carcasses were unearthed during excavation activities, the drums were set aside for fingerprint analyses in order to determine chemical compatability. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Monday, December 28, 1998

FOR DATES: From 7:00 a.m. Sunday, Dec. 27, 1998 to 7:00 a.m. Monday, Dec. 28, 1998.

WEATHER: Cloudy. The temperature range was 12oF to 37oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 792.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 33.2 tons per hour while on line. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, excavation and feed preparation. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-12-21-98-310). Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Danny Zion for William Werntges (USACE)


Sunday, December 27, 1998

FOR DATES: From 7:00 a.m. Saturday, Dec. 26, 1998 to 7:00 a.m. Sunday, Dec. 27, 1998.

WEATHER: Clear. The temperature range was 12oF to 39oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 706.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.7 tons per hour while on line. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations and feed preparation. Excavation on the south face of the exclusion zone was performed on this date. No unusual material or debris was noted in the excavated area. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-12-22-98-311) for bin 6. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, December 26, 1998

FOR DATES: From 7:00 a.m. Friday, Dec. 25, 1998 to 7:00 a.m. Saturday, Dec. 26, 1998.

WEATHER: Clear. The temperature range was 12oF to 31oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 643.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 26.9 tons per hour while on line. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations and feed preparation. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-12-20-98-309) for bin 1. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Greg Beauseigneur for William Werntges (USACE)


Friday, December 25, 1998

FOR DATES: From 7:00 a.m. Thursday, Dec. 24, 1998 to 7:00 a.m. Friday, Dec. 25, 1998.

WEATHER: Partly cloudy. The temperature range was 12oF to 32oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 592.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 24.8 tons per hour while on line. Related ongoing site activities include feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP)operations, feed preparation, debris handling, debris separation, lime mixing, ash handling and ash backfilling. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program. Received fly ash analytical report (fly-ash-12-19-98-062) for the contents of bin A. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of chromium, lead and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Thursday, December 24, 1998

FOR DATES: From 7:00 a.m. Wednesday, Dec. 23, 1998 to 7:00 a.m. Thursday, Dec. 24, 1998.

WEATHER: Partly cloudy. The temperature range was 15oF to 22oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 23.9 hours and 654.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.4 tons per hour while on line. Related ongoing site activities include debris separation and feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, debris separation, debris handling, ash handling and ash backfilling. A survey of the excavation bottom was performed. Excavation activities were continued near the location of the former debris separation building location. The lime silo was activated and filled on this date. The contractor continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-12-17-98-306, bot-ash-12-18-98-307 and bot-ash-12-19-98-308) for bin 4, bin 3 and bin 2 respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Wednesday, December 23, 1998

FOR DATES: From 7:00 a.m. Tuesday, Dec. 22, 1998 to 7:00 a.m. Wednesday, Dec. 23, 1998.

WEATHER: Windy and cold. The temperature range was 15oF to 39oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 23.8 hours and 716.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.1 tons per hour while on line. Related ongoing site activities include debris separation and feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, debris separation, debris handling, ash handling and ash backfilling. The metal framework of the debris building sprung structure was decontaminated and moved to the support zone. A survey of the excavation bottom was performed. Excavation activities were continued southeast of the former debris separation building location. The contractor continued with the real time perimeter air monitoring (PAM) program. One roll-off of non-hazardous metal debris and PPE was shipped off-site to a hazardous waste landfill.

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Sandra Downs, Larry Newcomer, Mike Welch (PADEP), Peter Shaw (USEPA)

PREPARED BY: Steve Gibson for Chad Thompson (USACE)


Tuesday, December 22, 1998

FOR DATES: From 7:00 a.m. Monday, Dec. 21, 1998 to 7:00 a.m. Tuesday, Dec. 22, 1998.

WEATHER: Overcast with rain. The temperature range was 40oF to 53oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 23.7 hours and 644.5 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.2 tons per hour while on line. Related ongoing site activities include debris separation, site housekeeping and feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, WWTP operations, feed preparation, debris separation, debris handling, ash handling and ash backfilling. Excavation activities were continued directly south of the former debris separation building location. The contractor continued with the real time perimeter air monitoring (PAM) program. Received the bottom ash analytical report (bot-ash-12-16-98-305) for bin 7. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for William Werntges (USACE)


Monday, December 21, 1998

FOR DATES: From 7:00 a.m. Sunday, Dec. 20, 1998 to 7:00 a.m. Monday, Dec. 21, 1998.

WEATHER: Overcast. The temperature range was 39oF to 42oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 16.25 hours and 433.9 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 26.7 tons per hour while on line. Downtime was due to a broken idler sprocket on the dry ash drag. Related ongoing site activities include debris separation, site housekeeping and feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, excavation, feed preparation, debris separation, debris handling, ash handling and ash backfilling. The contractor continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for William Werntges (USACE)


Sunday, December 20, 1998

FOR DATES: From 7:00 a.m. Saturday, Dec. 19, 1998 to 7:00 a.m. Sunday, Dec. 20, 1998.

WEATHER: Overcast with some rain. The temperature range was 32oF to 41oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 22.6 hours and 609.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.0 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, ash handling, and ash backfilling. The contractor continued to dismantle the debris separation building. Excavation activities continued along the southern face of the exclusion zone directly south of the debris separation building. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for William Werntges (USACE)


Saturday, December 19, 1998

FOR DATES: From 7:00 a.m. Friday, Dec. 18, 1998 to 7:00 a.m. Saturday, Dec. 19, 1998.

WEATHER: Clear and sunny. The temperature range was 29oF to 39oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 760 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 31.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. The contractor continued to dismantle the debris separation building. Excavation activities were continued along the southern face of the exclusion zone directly south of the debris separation building. Continued with the real time perimeter air monitoring (PAM) program. AirRecon completed the third quarterly stacktesting event for metals and particulate.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Kurt Davis (AIR), Sandra Downs (PADEP).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Friday, December 18, 1998

FOR DATES: From 7:00 a.m. Thursday, Dec. 17, 1998 to 7:00 a.m. Friday, Dec. 18, 1998.

WEATHER: Overcast with light snow. The temperature range was 30oF to 41oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 23.9 hours and 703.4 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 29.4 tons per hour while on line. Related ongoing site activities include debris separation, site housekeeping and feed preparation. The third and final run of metals and particulate emissions stack testing for the third quarter of the operations phase of the project is scheduled to continue during the morning hours of Friday, December 18, 1998.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, UST handling, feed preparation, debris separation, debris handling, wood shredding, rock cleaning, rock backfilling, ash handling and ash backfilling. Operations crews continued to dismantle the debris separation building as part of the close out plan for the exclusion zone. No excavation was performed on this date. The contractor continued with the real time perimeter air monitoring (PAM) program. AirRecon, the stack sampling contractor, performed the 1st and 2nd run of the quarterly stack tests. Received bottom ash analytical report (bot-ash-12-13-98-302) for the contents of bin 1. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: None.

V. NOTED VISITORS: Chris McBride (Focus Environmental), Kurt Davis (AIR).

PREPARED BY: Steve Gibson for William Werntges (USACE)


Thursday, December 17, 1998

FOR DATES: From 7:00 a.m. Wednesday, Dec. 16, 1998 to 7:00 a.m. Thursday, Dec. 17, 1998.

WEATHER: Overcast. The temperature range was 22oF to 37oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 17.9 hours and 513.7 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is attributed to the stoppage of feed in order to repair a gap that had developed between two of the floor plates of the Dry Ash Drag (DAD) conveyor. The TDF processed contaminated soil at an average rate of 28.7 tons per hour while on line. Related ongoing site activities include debris separation, site housekeeping and feed preparation. Metals and particulate emissions stack testing for the third quarter of the operations phase of the project is scheduled to begin during the morning hours of Thursday, December 17, 1998.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, drum handling, feed preparation, debris separation, debris handling, wood shredding, rock cleaning, rock backfilling, ash handling and ash backfilling. Operations crews began dismantling the debris separation building as part of the close out plan for the exclusion zone. Excavation was continued directly south of the debris separation building. A open-ended storage tank carcass was unearthed. The tank was filled with soil and no real-time VOC readings were recorded near the tank. The contractor continued with the real time perimeter air monitoring (PAM) program. AirRecon, the stack testing contractor, mobilized to the site in order to begin stack testing on Thursday the 17th of December.

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination.

IV. PROBLEMS NOTED OR ANTICIPATED: As stated above, hazardous waste feed was stopped at 1105 hours in order to purge the kiln and perform maintenance to the DAD cooling plates. A gap had developed between the first two floor plates of the DAD. Maintenance was performed in order to seal the gap and keep dry ash from leaking onto the TDF pad. Hazardous waste feed resumed at 1708 hours.

V. NOTED VISITORS: None.

PREPARED BY: Steve Gibson for William Werntges (USACE)


Wednesday, December 16, 1998

FOR DATES: From 7:00 a.m. Tuesday, Dec. 15, 1998 to 7:00 a.m. Wednesday, Dec. 16, 1998.

WEATHER: Clear and cold. The temperature range was 20oF to 52oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 23.9 hours and 651.1 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.2 tons per hour while on line. Related ongoing site activities include debris separation, site housekeeping and feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris separation, debris handling, wood shredding, rock cleaning, rock backfilling, ash handling and ash backfilling. Operations crews began dismantling the debris separation building as part of the close out plan for the exclusion zone. Minor excavation activities were performed adjacent to the debris separation building. The contractor continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulate emissions is scheduled for Thursday and Friday of this week, December 17th and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandra Downs and Larry Newcomer (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, December 15, 1998

FOR DATES: From 7:00 a.m. Monday, Dec. 14, 1998 to 7:00 a.m. Tuesday, Dec. 15, 1998.

WEATHER: Clear and cold. The temperature range was 21oF to 42oF.

I. CURRENT SITUATION: The contractor, OHM, is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was online for 17.7 hours and 433.2 tons of contaminated soils were thermally treated during the time frame of this report. Off-line time is primarily attributed to the stoppage of hazardous waste feed in order to perform repairs to the Wet Ash Drag's drive chain and tail shaft. The TDF processed contaminated soil at an average rate of 24.5 tons per hour while on line. Related ongoing site activities include debris separation, site housekeeping and feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, debris separation, debris handling and ash handling. Excavation activities were performed adjacent to the debris separation building. The contractor continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulate emissions is scheduled for Thursday and Friday of this week, December 17th and 18th.

IV. PROBLEMS NOTED OR ANTICIPATED: The TDF was off line for approximately 6 hours to repair the wet ash drag.

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, December 14, 1998

FOR DATES: From 7:00 a.m. Sunday, Dec. 13, 1998 to 7:00 a.m. Monday, Dec. 14, 1998.

WEATHER: Clear and cold. The temperature range was 25oF to 42oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) has resumed thermal treatment of contaminated soil. The contractor, OHM, restarted waste feed at 0903 hours on Sunday after a 7 day maintenance shutdown. The TDF was online for 21.72 hours and 594.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 27.4 tons per hour while on line. Related ongoing site activities include TDF maintenance, plant winterization, site housekeeping and feed preparation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, site housekeeping, ash backfill, debris handling and real time perimeter air monitoring (PAM) program. No bottom ash or fly ash analytical reports were received during this reports coverage.

III. FUTURE PLANS: Hazardous waste feed resumed on Sunday, December 13, 1998. Plans for this week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulate emissions is scheduled for Thursday and Friday of this week.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: John Ricard for David Modricker (USACE)


Sunday, December 13, 1998

FOR DATES: From 7:00 a.m. Saturday, Dec. 12, 1998 to 7:00 a.m. Sunday, Dec. 13, 1998.

WEATHER: Clear and breezy. The temperature range was 23oF to 39oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently heating up to temperature in order to resume the thermal treatment of contaminated soil. The "hot hold" temperatures were reached at approximately 6:00 p.m. on 12-12-98. The monthly interlock test checks were performed and observed by QA/QC staff. On Saturday, at 0842 hrs, an unplanned thermal relief valve (TRV) event occurred when a pressurized air line froze. The air line provides air to a pneumatic cylinder that keeps the TRV closed. The frozen air line was repaired and the TRV closed at 0940 hrs. No waste feed was being processed at that time due to the current status and all required notifications were made. The contractor, OHM, continues to work 24-hours a day, 7 days a week maintaining the TDF. Related ongoing site activities include site housekeeping, feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued maintenance to the TDF. Major maintenance activities included removing the build up of fly ash in the quench tower, maintenance and inspection of the rotary kiln, maintenance to the Dry Ash Drag floor plates and the reinstallation of the new dual feed screws. Site activities included debris separation, feed preparation, site housekeeping and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

No bottom ash or fly ash analytical reports were received during this reports coverage.

III. FUTURE PLANS: Hazardous waste feed is tentatively scheduled to resume in the early morning hours on Sunday, December 13, 1998. Plans for next week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulate emissions is scheduled for next week, once feed resumes.

IV. PROBLEMS NOTED OR ANTICIPATED: Reference section 1 above for a description of an unplanned TRV on Saturday, December 12, 1998. Waste feed was off (had not resumed) prior to the TRV opening.

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, December 12, 1998

FOR DATES: From 7:00 a.m. Friday, Dec. 11, 1998 to 7:00 a.m. Saturday, Dec. 12, 1998.

WEATHER: Clear and breezy. The temperature range was 24oF to 44oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently heating up to perform the monthly interlock checks and resume the thermal treatment of contaminated soil. Burners were lit during the period of this report including one kiln burner at 0030 hours, the first Secondary Combustion Chamber (SCC) burner at 0306 hours and the second SCC burner at 0522 hours on Friday, December 11, 1998. The contractor, OHM, continues to work 24-hours a day, 7 days a week maintaining the TDF. Related ongoing site activities include site housekeeping, feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued maintenance to the TDF. Major maintenance activities included removing the build up of fly ash in the quench tower, maintenance and inspection of the rotary kiln, maintenance to the Dry Ash Drag floor plates and the reinstallation of the new dual feed screws. Site activities included Waste Water Treatment Plant (WWTP) operations, debris separation, feed preparation, site housekeeping and ash backfilling. Excavation activities were limited to the removal of soil sediment from sediment basin 2. Continued with the real time perimeter air monitoring (PAM) program.

Received bottom ash analytical reports (bot-ash-12-04-98-299, bot-ash-12-05-98-300, bot-ash-12-06-98-301) for bin 4, bin 3 and bin 2 respectively. Test results indicated that the thermally treated soil sampled from bin 3 and bin 2 (bot-ash-12-05-98-300 and bot-ash-12-06-98-301) met the project requirements for unrestricted backfilling on-site. Test results indicated that the contents of bin 4 (bot-ash-12-04-98-299) met the project requirement for restricted backfilling on-site (i.e. above elevation 553) due to the concentration of arsenic being greater than the PADEP drinking water standard, but less than 25 times the standard. Received fly ash analytical report (fly-ash-12-07-98-061) for bin E. Test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site due to the concentrations of arsenic, cadmium, chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: There will be a monthly interlock check performed on Saturday, December 12, 1998 in the late evening hours. The interlock checks will be performed in order to test the TDF's automatic shut down system. Hazardous waste feed is tentatively scheduled to resume in the early morning hours on Sunday, December 13, 1998. Plans for next week include the continuation of excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Stack testing for metals and particulate emissions is scheduled for next week, once feed resumes.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, December 11, 1998

FOR DATES: From 7:00 a.m. Thursday, Dec. 10, 1998 to 7:00 a.m. Friday, Dec. 11, 1998.

WEATHER: Clear. The temperature range was 23oF to 45oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line in order to perform maintenance on the internal and external components of the incinerator. The contractor, OHM, continues to work 24-hours a day, 7 days a week maintaining the TDF. Related ongoing site activities include site housekeeping, feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued maintenance to the TDF. Major maintenance activities included removing built up fly ash from quench tower and maintenance to kiln, wet ash drag, scrubber and generator. Preparations are being made to reinstall the feed screws. Sludge was removed from the North end of the sedimentation basin. Continued trommel setup in the feed preparation building. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Preventative maintenance on the external and internal components of the incinerator will continue and the feed screws will be reinstalled. During the TDF down time, the Contractor will continue excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Hazardous waste feed is scheduled to resume this weekend or early next week. Stack testing for metals and particulate emissions is scheduled next week after feed resumes.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandra Downs, PADEP

PREPARED BY: Danny Zion for David Modricker (USACE)


Thursday, December 10, 1998

FOR DATES: From 7:00 a.m. Wednesday, Dec. 9, 1998 to 7:00 a.m. Thursday, Dec. 10, 1998.

WEATHER: Clear. The temperature range was 24oF to 48oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line in order to perform maintenance on the internal and external components of the incinerator. The contractor, OHM, continues to work 24-hours a day, 7 days a week maintaining the TDF. Related ongoing site activities include site housekeeping, feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued maintenance to the TDF. Major maintenance activities included acid washing the Dry Ash Drag floor plate cooling system, replacement of kiln breech seals and kiln burners, and miscellaneous TDF maintenance. The contractor also continued Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, rock cleaning and rock backfilling. No new excavation was performed on this date. A trommel was moved from the debris separation building to the feed preparation building. Continued with the real time perimeter air monitoring (PAM) program.

III FUTURE PLANS: Preventative maintenance on the external and internal components of the incinerator and replacement of the current feed screws will continue through the week. During the TDF down time, the Contractor will continue excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Hazardous waste feed is scheduled to resume late this week or early next week. Stack testing for metals and particulate emissions is scheduled after feed resumes during the third week in December, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE)


Wednesday, December 9, 1998

FOR DATES: From 7:00 a.m. Tuesday, Dec. 8, 1998 to 7:00 a.m. Wednesday, Dec. 9, 1998.

WEATHER: Overcast with rain. The temperature range was 32oF to 46oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line in order to perform maintenance on the internal and external components of the incinerator. The contractor, OHM, continues to work 24-hours a day, 7 days a week maintaining the TDF. Related ongoing site activities include site housekeeping, feed preparation and debris separation. There was a planned Thermal Relief Valve (TRV) opening initiated on Wednesday, December 9, 1998 at 0650 hours. The TRV was opened in order to perform additional maintenance to the internal components of the incinerator. Feed to the incinerator had been stopped for over 48 hours prior to the opening. The necessary authorities were notified before the TRV opening was initiated.

II. ACTIONS TAKEN: The contractor continued maintenance to the TDF. Major maintenance activities included the replacement of the Wet Ash Drag drive chain, inspection and cleaning of the cyclone, maintenance to the weighbelt conveyance system and maintenance to the Dry Ash Drag floor plate cooling system. The contractor also continued site house keeping, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris separation, rock cleaning and rock backfilling. No new excavation was performed on this date. The monthly survey of the exclusion zone was performed. Continued with the real time perimeter air monitoring (PAM) program and began the weekly OSHA method 93 24-hour sampling event for BNA at the on-site and off-site air monitoring stations.

III FUTURE PLANS: Preventative maintenance on the external and internal components of the incinerator and replacement of the current feed screws will continue through the week. During the TDF down time, the Contractor will continue excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Hazardous waste feed is scheduled to resume late this week or early next week. Stack testing for metals and particulate emissions is scheduled after feed resumes during the third week in December, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Steve Gibson for David Modricker (USACE)


Tuesday, December 8, 1998

FOR DATES: From 7:00 a.m. Monday, Dec. 7, 1998 to 7:00 a.m. Tuesday, Dec. 8, 1998.

WEATHER: Overcast. The temperature range was 42oF to 67oF.

I. CURRENT SITUATION: The Thermal Destruction Facility (TDF) is currently off-line in order to perform maintenance on the internal and external components of the incinerator. The contractor, OHM, continues to work 24-hours a day, 7 days a week maintaining the TDF. Related ongoing site activities include site housekeeping, feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued maintenance to the TDF. Major maintenance activities included the removal of the dual feed screws, inspection and maintenance to the ash drag conveyors and maintenance to the weighbelt. The contractor also continued Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, ash handling and ash backfilling. Debris separation activities are currently being performed 24-hours a day in order to process, decontaminate and backfill site rock. Excavation activities were continued along the southern face of the exclusion zone. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-12-01-98-296, bot-ash-12-02-98-297 and bot-ash-12-03-98-298) for bin 7, bin 6 and bin 5, respectively. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-12-02-98-060) for bin D. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III FUTURE PLANS: Preventative maintenance on the external and internal components of the incinerator and replacement of the current feed screws will continue through the week. A thermal relief valve (TRV) opening is planned for Wednesday, December 10, 1998. The proper authorities have been notified of the TRV opening. During the TDF down time, the Contractor will continue excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Hazardous waste feed is scheduled to resume late this week or early next week. Stack testing for metals and particulate emissions is scheduled after feed resumes during the third week in December, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Kurt Davis (AIR)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Monday, December 7, 1998

FOR DATES: From 7:00 a.m. Sunday, Dec. 6, 1998 to 7:00 a.m. Monday, Dec. 7, 1998.

WEATHER: Partly cloudy and mild. The temperature range was 49oF to 62oF.

I. CURRENT SITUATION: Hazardous waste feed to the Thermal Destruction Facility (TDF) was stopped at 3:02 a.m. today in preparation of the planned preventative maintenance shutdown. The TDF was on-line for 19.9 hours and 696.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued feed preparation, lime mixing, rock cleaning, rock backfilling, ash handling, ash backfilling, and continued the real time perimeter air monitoring (PAM) program during the period of this report. Received ash analytical report (bot-ash-12-01-98-296). The ash met project requirements for unrestricted backfill on site.

III FUTURE PLANS: Preventative maintenance on the external and internal components of the incinerator and replacement of the current feed screws will continue through the week. A thermal relief valve (TRV) opening is planned in the next couple days as part of the maintenance work. Proper authorities will be notified prior to the TRV opening. During the TDF down time, the Contractor will continue excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. Hazardous waste feed is scheduled to resume the week of December 13, 1998. Stack testing for metals and particulate emissions is scheduled after feed resumes during the third week in December, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE)


Sunday, December 6, 1998

FOR DATES: From 7:00 a.m. Saturday, Dec. 5, 1998 to 7:00 a.m. Sunday, Dec. 6, 1998.

WEATHER: Partly cloudy and mild. The temperature range was 47oF to 59oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 849.3 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, rock cleaning, rock backfilling, ash handling, and ash backfilling. Continued with the real time perimeter air monitoring (PAM) program.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a cold shutdown of the TDF scheduled to begin the early hours of December 7, 1998. The contractor will perform preventative maintenance on the external and internal components of the incinerator and replace the current feed screws. Hazardous waste feed is scheduled to resume early in the week of December 13, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Richard Lupold for David Modricker (USACE)


Saturday, December 5, 1998

FOR DATES: From 7:00 a.m. Friday, Dec. 4, 1998 to 7:00 a.m. Saturday, Dec. 5, 1998.

WEATHER: Partly cloudy and mild with light evening rain. The temperature range was 51oF to 68oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 832.2 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 34.8 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, ash backfilling and an excavation bottom survey. Excavation activities were continued along the southern face of the exclusion zone adjacent to the debris separation building. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-11-29-98-294 and bot-ash-11-30-98-295) for bin 2 and bin 1 respectively. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There is a cold shutdown of the TDF scheduled to begin the early hours of December 7, 1998. The contractor will perform preventative maintenance on the external and internal components of the incinerator and replace the current feed screws. Hazardous waste feed is scheduled to resume early in the week of December 13, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Steve Gibson for David Modricker (USACE)


Friday, December 4, 1998

FOR DATES: From 7:00 a.m. Thursday, Dec. 3, 1998 to 7:00 a.m. Friday, Dec. 4, 1998.

WEATHER: Warm & partly cloudy. The temperature range was 37oF to 66oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 726.7 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued directly south of the debris separation building. Continued with the real time perimeter air monitoring (PAM) program, and began the modified OSHA method 93 air sampling event for BNA at the Off-site Air Monitoring (OAM) and PAM stations.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There will be a cold shutdown of the TDF starting December 7, 1998 in order to perform preventative maintenance on the external and internal components of the incinerator and to replace the feed screws. Stack testing is scheduled for the third week in December, 1998.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Chad Thompson for David Modricker (USACE)


Thursday, December 3, 1998

FOR DATES: From 7:00 a.m. Wednesday, Dec. 2, 1998 to 7:00 a.m. Thursday, Dec. 3, 1998.

WEATHER: Clear, breezy and mild. The temperature range was 28oF to 66oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.9 hours and 732 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 30.6 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, Waste Water Treatment Plant (WWTP) operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued directly south of the debris separation building. Continued with the real time perimeter air monitoring (PAM) program, and began the modified OSHA method 93 air sampling event for BNA at the Off-site Air Monitoring (OAM) and PAM stations.

Received bottom ash analytical report (bot-ash-11-28-98-293) for bin 3. Test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site. Received fly ash analytical report (fly-ash-11-28-98-059) for bin C. Analytical test results indicated that the fly ash materials met the project requirements for restricted backfilling on-site (i.e. above elevation 553) due to the concentrations of cadmium, chromium and selenium being greater than the PADEP drinking water standard, but less than 25 times the standard.

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There will be a cold shutdown of the TDF starting December 7, 1998 in order to perform preventative maintenance on the external and internal components of the incinerator and to replace the feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Chad Thompson for David Modricker (USACE)


Wednesday, December 2, 1998

FOR DATES: From 7:00 a.m. Tuesday, Dec. 1, 1998 to 7:00 a.m. Wednesday, Dec. 2, 1998.

WEATHER: Clear and mild. The temperature range was 29oF to 54oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.7 hours and 851.6 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.9 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, rock cleaning, rock backfilling, ash handling, and ash backfilling. Excavation activities were continued directly south of the debris separation building. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical report (bot-ash-11-27-98-292) for bin 4. Analytical test results indicated that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There will be a cold shutdown of the TDF starting December 7, 1998 in order to perform preventative maintenance on the external and internal components of the incinerator and to replace the feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: Sandy Downs (PADEP)

PREPARED BY: Kenneth R. Hilton for David Modricker (USACE)


Tuesday, December 1, 1998

FOR DATES: From 7:00 a.m. Monday, Nov. 30, 1998 to 7:00 a.m. Tuesday, Dec. 1, 1998.

WEATHER: Partly cloudy and mild. The temperature range was 40oF to 63oF.

I. CURRENT SITUATION: OHM is working 24 hours a day, seven days a week operating the Thermal Destruction Facility (TDF). The TDF was on-line for 23.8 hours and 844.8 tons of contaminated soils were thermally treated during the time frame of this report. The TDF processed contaminated soil at an average rate of 35.5 tons per hour while on line. Related ongoing site activities include feed preparation and debris separation.

II. ACTIONS TAKEN: The contractor continued TDF operations, feed preparation, lime mixing, debris handling, debris separation, wood shredding, ash handling, and ash backfilling. No new excavation was performed on this date. Continued with the real time perimeter air monitoring (PAM) program. Received bottom ash analytical reports (bot-ash-11-25-98-290 and bot-ash-11-26-98-291) for bins 6 and 5, respectively. Analytical test results indicate that the thermally treated soil met the project requirements for unrestricted backfilling on-site (i.e. above elevation 547).

III. FUTURE PLANS: Major activities for the next week include operation of the TDF, excavation, drum handling (if encountered), feed preparation, debris handling, ash handling, ash backfilling, off-site & perimeter air monitoring, WWTP operations, and rock cleaning & inorganic debris decontamination. There will be a cold shutdown of the TDF starting December 7th 1998 in order to perform preventative maintenance on the external and internal components of the incinerator and to replace the feed screws.

IV. PROBLEMS NOTED OR ANTICIPATED: None

V. NOTED VISITORS: None

PREPARED BY: Danny Zion for David Modricker (USACE)

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.